A charge of +3µC is distributed along a rod which on the x-axis extending from x=1m to

4m. a) What is the linear charge density of the rod? b) Find the electric field at the origin due to the charged rod.


1
Expert's answer
2021-04-26T17:12:25-0400

Answer

a. Linear charge density of rod

"\lambda=\frac{Q}{l}\\=\frac{3\times10^{-6}}{4-1}\\=1\mu C/m"

b. Electric field is given at origin

"E=\frac{2k\lambda}{d(d+l)}\\=\frac{2\times9\times10^9\times10^{-6}}{1\times4}\\=4500N/C."